OpenStudy (anonymous):

The volumes of two similar solids are 1000 m and 343 m. The surface area of the larger one is 500 m. What is the surface area of the smaller one?

OpenStudy (anonymous):

Is this a multiple choice question or essay?

OpenStudy (anonymous):

a) 172m. B) 1715 m. C) 245 m. D) none of these These are the choices for answer

OpenStudy (b87lar):

Suppose the solid has one parameter (radius). Volume is a cubic function of that. The larger volume is \[10^3=1000\] the smaller volume is \[7^3=343\]

OpenStudy (b87lar):

Surface is a quadratic function of the same parameter.

OpenStudy (anonymous):

I don't understand what you mean. Can you show me equation to figure this?

OpenStudy (b87lar):

the ratio of the volumes is \[\frac{343}{1000}=(\frac{7}{10})^3\] so the ratio of the surfaces will be \[(\frac{7}{10})^2=\frac{49}{100}\]

OpenStudy (b87lar):

now we know the larger surface is 500

OpenStudy (b87lar):

with that and the ratio we can get the smaller surface by multiplying \[\frac{49}{100} 500\]

OpenStudy (b87lar):

Does it make sense?

OpenStudy (anonymous):

Yes, thank you. That ratio helps so much! I get it now

OpenStudy (b87lar):

my pleasure